OpenStudy (anonymous):

Triangle ABC has sides with lengths 14, 17, and 26. Is the triangle obtuse, acute, or a right triangle?

OpenStudy (anonymous):

do a^2+b^2=c^2

OpenStudy (anonymous):

it is obtuse

OpenStudy (anonymous):

If c > b > a: c² = a² + b² => right triangle c² > a² + b² => obtuse triangle c² < a² + b² => acute triangle 26² = 676 14² + 17² = 485 676 > 485 => ABC is an obtuse triangle ...